OpenStudy (anonymous):

Find the points on the graph where the tangent is horizontal or vertical (dy/dx=(dy/dt)/(dx/dt)) x=t^3-3t^2, y=2t^2+8t-5

OpenStudy (campbell_st):

differentiate both functions with respect to t. \[\frac{dx}{dt} = 3t^2 - 6t\] \[\frac{dy}{dt} = 4t + 8\] now using the hint \[\frac{dy}{dx} = \frac{4t + 8}{3t^2 - 6t}\] you may need to check if it can be simplified. where is the gradient of the tangent horizontal... when the 0 = 4t + 8 when is it vertical when 3t^2 - 6t = 0 hope this helps

OpenStudy (anonymous):

thank you for your clear reply! i followed your steps to get that the solutions for t are -2 (horizontal); 0,2 (vertical) how do i check to make sure that the solutions are "real" or is this the end of the problem?

OpenStudy (campbell_st):

this is the end... the solutions to the vertical are asymptotes as you can't divide by zero...

OpenStudy (anonymous):

ok thank you :)